# Euclid Quick Data Release (Q1)

### v1

@@
Euclid spectroscopy of QSOs. 1. Identification and redshift determination of 3500 bright QSOs
@@
Euclid Collaboration: Y. Fu    R. Bouwens    K. I. Caputi    D. Vergani    M. Scialpi    B. Margalef-Bentabol    L. Wang    M. Bolzonella    M. Banerji    E. Bañados    A. Feltre    Y. Toba    J. Calhau    F. Tarsitano    P. A. C. Cunha    A. Humphrey    G. Vietri    F. Mannucci    S. Bisogni    F. Ricci    H. Landt    L. Spinoglio    T. Matamoro Zatarain    D. Stern    M. J. Page    D. M. Alexander    G. Zamorani    W. Roster    M. Salvato    Y. Copin    J. G. Sorce    D. Scott    Y.-H. Zhang    E. Lusso    J. Wolf    D. Yang    H. J. A. Rottgering    B. Laloux    M. Siudek    S. Belladitta    Q. Liu    V. Allevato    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    A. Balestra    S. Bardelli    P. Battaglia    A. Biviano    E. Branchini    M. Brescia    J. Brinchmann    S. Camera    G. Cañas-Herrera    V. Capobianco    C. Carbone    J. Carretero    S. Casas    M. Castellano    G. Castignani    S. Cavuoti    K. C. Chambers    A. Cimatti    C. Colodro-Conde    G. Congedo    C. J. Conselice    L. Conversi    A. Costille    F. Courbin    H. M. Courtois    M. Cropper    A. Da Silva    H. Degaudenzi    G. De Lucia    C. Dolding    H. Dole    F. Dubath    C. A. J. Duncan    X. Dupac    S. Dusini    S. Escoffier    M. Fabricius    M. Farina    R. Farinelli    S. Ferriol    F. Finelli    P. Fosalba    N. Fourmanoit    M. Frailis    E. Franceschi    P. Franzetti    M. Fumana    S. Galeotta    K. George    W. Gillard    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    H. Hoekstra    W. Holmes    I. M. Hook    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    A. Kiessling    B. Kubik    M. Kümmel    M. Kunz    H. Kurki-Suonio    R. Laureijs    A. M. C. Le Brun    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    G. Mainetti    D. Maino    E. Maiorano    O. Mansutti    S. Marcin    O. Marggraf    K. Markovic    M. Martinelli    N. Martinet    F. Marulli    R. J. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier    M. Meneghetti    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    R. Nakajima    C. Neissner    R. C. Nichol    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    W. J. Percival    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    R. Rebolo    A. Renzi    J. Rhodes    G. Riccio    E. Romelli    M. Roncarelli    E. Rossetti    R. Saglia    Z. Sakr    D. Sapone    B. Sartoris    M. Schirmer    P. Schneider    T. Schrabback    M. Scodeggio    A. Secroun    E. Sefusatti    G. Seidel    S. Serrano    P. Simon    C. Sirignano    G. Sirri    L. Stanco    J.-L. Starck    J. Steinwagner    C. Surace    P. Tallada-Crespí    D. Tavagnacco    A. N. Taylor    H. I. Teplitz    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    L. Valenziano    J. Valiviita    T. Vassallo    A. Veropalumbo    D. Vibert    Y. Wang    J. Weller    A. Zacchei    E. Zucca    M. Ballardini    E. Bozzo    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    L. Gabarra    W. G. Hartley    M. Huertas-Company    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. A. Nucita    A. Pezzotta    M. Pöntinen    C. Porciani    I. Risso    V. Scottez    M. Sereno    M. Tenti    M. Viel    M. Wiesmann    Y. Akrami    S. Alvi    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    E. Aubourg    D. Bertacca    M. Bethermin    L. Bisigello    A. Blanchard    L. Blot    M. Bonici    S. Borgani    M. L. Brown    S. Bruton    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    T. Castro    F. Cogato    S. Conseil    A. R. Cooray    O. Cucciati    G. Daste    F. De Paolis    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    P. Dimauro    P.-A. Duc    M. Y. Elkhashab    A. Enia    Y. Fang    A. G. Ferrari    A. Finoguenov    F. Fontanot    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    E. Gaztanaga    F. Giacomini    F. Gianotti    G. Gozaliasl    M. Gray    M. Guidi    C. M. Gutierrez    A. Hall    C. Hernández-Monteagudo    H. Hildebrandt    J. Hjorth    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    J. Kim    C. C. Kirkpatrick    S. Kruk    V. Le Brun    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    T. I. Liaudat    A. Loureiro    J. Macias-Perez    M. Magliocchetti    C. Mancini    R. Maoli    C. J. A. P. Martins    L. Maurin    M. Miluzio    P. Monaco    C. Moretti    G. Morgante    S. Nadathur    K. Naidoo    P. Natoli    A. Navarro-Alsina    S. Nesseris    D. Paoletti    F. Passalacqua    K. Paterson    L. Patrizii    A. Pisani    D. Potter    S. Quai    M. Radovich    P.-F. Rocci    G. Rodighiero    S. Sacquegna    M. Sahlén    D. B. Sanders    E. Sarpa    C. Scarlata    A. Schneider    D. Sciotti    E. Sellentin    F. Shankar    L. C. Smith    E. Soubrie    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    G. Verza    P. Vielzeuf    A. Viitanen    N. A. Walton    J. R. Weaver
@@
The slitless spectroscopy mode of the Near-Infrared Spectrometer and Photometer (NISP) onboard the Euclid telescope has enabled efficient spectroscopy of objects within a large field of view. Nevertheless, the relatively low spectral resolution, overlapping spectra, and contamination pose challenges to source classification and redshift determination using the NISP spectra alone. In this work, we present a large and homogeneous sample of bright quasars identified from the Euclid Quick Data Release (Q1), constructed by combining high-purity candidate selections from Gaia and WISE with the new spectroscopic capabilities of Euclid. Through visual inspection of the Euclid spectra of these quasar candidates, we identify approximately 3500 quasars with reliable redshifts in the range of $0<z\lesssim 4.8$. We generate the first Euclid composite spectrum of quasars covering rest-frame near-ultraviolet (NUV) to near-infrared (NIR) wavelengths without telluric lines, which will be pivotal to NIR quasar spectral analysis. We obtain an empirical spectroscopic depth of $J_{\scriptscriptstyle\rm E}\lesssim 21.5$ and $H_{\scriptscriptstyle\rm E}\lesssim 21.3$ at the sensitivity of the Wide Field Survey, beyond which the number of securely identified quasars declines sharply. Accordingly, the sample presented in this paper comprises spectroscopically confirmed quasars brighter than these limits. We analyse VIS morphological parameters using Srsic and model-independent (CAS) metrics, and a deep-learning point spread function (PSF) fraction to track nuclear dominance. VIS morphologies show a clear redshift dependence: at low redshift ($z<0.5$), obvious host structures are common and a single Srsic model fits about half of the sources; at intermediate redshift ($0.5<z<2$), the nuclear component dominates, with 90% of the Srsic fits saturating at the upper index limit. In this intermediate redshift regime, $f_{{\scriptscriptstyle\rm PSF}}$ is available, and we use it as a more reliable compactness measure than the single-Srsic and CAS parameters to quantify nuclear versus host emission. We also explore the novel Euclid NIR colour space and discuss the role of these quasars in refining active galactic nuclei (AGN) selection techniques for future Euclid data releases. Our results highlight the potential of Euclid spectroscopy to advance quasar surveys and enable the construction of more complete AGN catalogues. The spectroscopic bright quasar catalogue of this work, and the composite quasar spectrum, will be available at https://cdsarc.cds.unistra.fr/.
@@ 1 Introduction
Powered by accreting supermassive black holes, quasars are among the most luminous and distant objects in the Universe. Quasars act as beacons that allow us to probe recent to early cosmic epochs and trace the largescale structure of the cosmos (e.g., 2017AJ....154...28B; 2020MNRAS.499..210N; 2023ARA&A..61..373F). Traditionally, quasar surveys have primarily relied on multi-colour selection techniques to isolate candidates, with subsequent slit or multi-fibre spectroscopy to confirm their nature and determine precise redshifts (e.g., 2002AJ....123.2945R; 2004MNRAS.349.1397C; 2015ApJS..221...27M; 2023ApJ...944..107C). Ground-based quasar surveys have adopted slitless spectroscopy to find quasars in an unbiased and cost-effective way. However, such slitless campaigns suffer from low spectral resolution, overlapping spectra, and contamination, often necessitating follow-up slit spectroscopy for confirmation (e.g., 1986ApJ...306..411S; 1991ApJS...75..273O; 1999AJ....117...40S).
@@
The advent of space-based observations and improvements in data reduction have substantially improved the quality of slitless spectroscopic data. For example, the Hubble Space Telescope’s (HST) Advanced Camera for Surveys (ACS) and Wide Field Camera 3 (WFC3) have produced high-quality grism data (e.g., Momcheva_2016ApJS_3dhst_survey; 2019ApJ...870..133E) that are processed with dedicated pipelines (e.g., 2009PASP..121...59K; 2019ascl.soft05001B). More recently, with the low background, high sensitivity, and high spatial resolution of the James Webb Space Telescope (JWST), the Near Infrared Imager and Slitless Spectrograph (NIRISS; 2022PASP..134b5002W; 2023PASP..135i8001D) onboard JWST has enabled unprecedented studies of distant galaxies (e.g., 2022ApJ...938L..13R; 2023ApJ...953...53S; 2024MNRAS.535.1067M).
@@
The European Space Agency’s (ESA) Euclid mission (EuclidSkyOverview) is designed to probe the dark matter and dark energy of the Universe by studying weak lensing and galaxy clustering over approximately one-third of the sky in both the optical and near-infrared using the Visible Camera (VIS; EuclidSkyVIS) and the Near-Infrared Spectrometer and Photometer (NISP; EuclidSkyNISP). A key feature of Euclid is the slitless spectroscopic mode of NISP, which can simultaneously capture spectra of sources in a large field of view of 0.57 $\deg^{2}$. Three red grisms covering the same ${\rm RG}_{\scriptscriptstyle\rm E}$ band (1206–1892 nm) are adopted in the Euclid Wide Survey to provide spectra with different dispersion directions of 0°, 180°, and 270° with respect to the detector columns. Dispersed slitless images of the three grims are combined to disentangle overlapping spectra from multiple sources, and generate the final clean spectra. The red grims have a resolving power of ${{\cal R}_{{\rm{R}}{{\rm{G}}_{\rm{E}}}}}>480$ for a source with a 0.5 diameter (EuclidSkyNISP; EuclidSkyOverview). This spectroscopic capability enables an unbiased census of bright quasars and enhances the efficiency of quasar discovery by leveraging the high sensitivity and spatial resolution from space. For example, 2025arXiv250613945B have recently discovered a $z=5.404$ quasar EUCL J181530.01+652054.0 using NISP spectroscopy.
@@
EP-Lusso provides a detailed prediction of the NISP spectroscopic mode for AGN using mock spectra. They demonstrate that redshift measurements are robust when the H $\alpha$ emission line is visible within the spectral coverage of ${\rm RG}_{\scriptscriptstyle\rm E}$ ($0.89<z<1.83$) at a line flux greater than $2\times 10^{-16}$ erg s-1 cm-2. Outside this redshift range, however, redshift measurements are inefficient due to low signal-to-noise ratio (S/N) or lack of prominent emission lines or both.
@@
Early investigations of AGN in the first Euclid Quick Data Release (Q1; Q1-TP001) already demonstrate the potential of Euclid for AGN science. For example, Q1-SP027 present an AGN candidate catalogue with a total of 229 779 objects selected with multiwavelength data, while Q1-SP003 identify Euclid counterparts to X-ray sources in the Deep Fields, most of which are AGN, using catalogues from eROSITA (2024A&A...682A..34M), XMM-Newton (2020A&A...641A.136W), and Chandra (2024ApJS..274...22E).
@@
In this work, we present a homogenous bright quasar sample identified with spectroscopic data of Q1, complementary to the efforts of Q1-SP027 and Q1-SP003. We select quasar candidates from all-sky optical and mid-infrared databases, and identify the objects based on prominent emission lines in the Euclid slitless spectra. Our analysis demonstrates that slitless spectroscopy, as implemented in the Euclid mission, not only overcomes some of the limitations of colour-based quasar selection techniques, but also provides a valuable data set for studying the large-scale structure of the Universe through quasar clustering and cross-correlation with galaxy and weak-lensing maps (e.g., 2003MNRAS.342..467M; 2015MNRAS.449.4326P; 2023ApJ...946...27P; 2023JCAP...11..043A). Spectral properties of these quasars obtained with multi-component fitting, including spectral indices, line widths, and black hole masses, will be reported in Euclid Collaboration: Calhau et al. (in preparation).
@@
The paper is organised as follows. Section 2 describes the sample selections of quasar candidates and the crossmatch to the Euclid spectral catalogue. Section 3 describes the identification and redshift determination procedure. We present our results in Sect. 4 and discuss their implications in Sect. 5. Finally, Sect. 6 concludes with a summary of our findings and prospects for future studies. All magnitudes are in the AB system (1983ApJ...266..713O) unless stated otherwise.

### v2

@@
Euclid spectroscopy of quasars. 1. Identification and redshift determination of 3500 bright quasarsthanks: The full catalogue of the identified quasars and the composite spectra are available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@
Euclid Collaboration: Y. Fu    R. Bouwens    K. I. Caputi    D. Vergani    M. Scialpi    B. Margalef-Bentabol    L. Wang    M. Bolzonella    M. Banerji    E. Bañados    A. Feltre    Y. Toba    J. Calhau    F. Tarsitano    P. A. C. Cunha    A. Humphrey    G. Vietri    F. Mannucci    S. Bisogni    F. Ricci    H. Landt    L. Spinoglio    T. Matamoro Zatarain    D. Stern    M. J. Page    D. M. Alexander    G. Zamorani    W. Roster    M. Salvato    Y. Copin    J. G. Sorce    D. Scott    Y.-H. Zhang    E. Lusso    J. Wolf    D. Yang    H. J. A. Rottgering    B. Laloux    M. Siudek    S. Belladitta    Q. Liu    V. Allevato    K. Kuijken    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    A. Balestra    S. Bardelli    P. Battaglia    A. Biviano    E. Branchini    M. Brescia    J. Brinchmann    S. Camera    G. Cañas-Herrera    V. Capobianco    C. Carbone    J. Carretero    S. Casas    M. Castellano    G. Castignani    S. Cavuoti    K. C. Chambers    A. Cimatti    C. Colodro-Conde    G. Congedo    C. J. Conselice    L. Conversi    A. Costille    F. Courbin    H. M. Courtois    M. Cropper    A. Da Silva    H. Degaudenzi    G. De Lucia    C. Dolding    H. Dole    F. Dubath    C. A. J. Duncan    X. Dupac    S. Dusini    S. Escoffier    M. Fabricius    M. Farina    R. Farinelli    S. Ferriol    F. Finelli    P. Fosalba    N. Fourmanoit    M. Frailis    E. Franceschi    P. Franzetti    M. Fumana    S. Galeotta    K. George    W. Gillard    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    H. Hoekstra    W. Holmes    I. M. Hook    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    A. Kiessling    B. Kubik    M. Kümmel    M. Kunz    H. Kurki-Suonio    R. Laureijs    A. M. C. Le Brun    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    G. Mainetti    D. Maino    E. Maiorano    O. Mansutti    S. Marcin    O. Marggraf    K. Markovic    M. Martinelli    N. Martinet    F. Marulli    R. J. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier Deceased    M. Meneghetti    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    R. Nakajima    C. Neissner    R. C. Nichol    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    W. J. Percival    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    R. Rebolo    A. Renzi    J. Rhodes    G. Riccio    E. Romelli    M. Roncarelli    E. Rossetti    R. Saglia    Z. Sakr    D. Sapone    B. Sartoris    M. Schirmer    P. Schneider    T. Schrabback    M. Scodeggio    A. Secroun    E. Sefusatti    G. Seidel    S. Serrano    P. Simon    C. Sirignano    G. Sirri    L. Stanco    J.-L. Starck    J. Steinwagner    C. Surace    P. Tallada-Crespí    D. Tavagnacco    A. N. Taylor    H. I. Teplitz    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    L. Valenziano    J. Valiviita    T. Vassallo    A. Veropalumbo    D. Vibert    Y. Wang    J. Weller    A. Zacchei    E. Zucca    M. Ballardini    E. Bozzo    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    L. Gabarra    W. G. Hartley    M. Huertas-Company    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. A. Nucita    A. Pezzotta    M. Pöntinen    C. Porciani    I. Risso    V. Scottez    M. Sereno    M. Tenti    M. Viel    M. Wiesmann    Y. Akrami    S. Alvi    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    E. Aubourg    D. Bertacca    M. Bethermin    L. Bisigello    A. Blanchard    L. Blot    M. Bonici    S. Borgani    M. L. Brown    S. Bruton    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    T. Castro    F. Cogato    S. Conseil    A. R. Cooray    O. Cucciati    G. Daste    F. De Paolis    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    P. Dimauro    P.-A. Duc    M. Y. Elkhashab    A. Enia    Y. Fang    A. G. Ferrari    A. Finoguenov    F. Fontanot    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    E. Gaztanaga    F. Giacomini    F. Gianotti    G. Gozaliasl    M. Gray    M. Guidi    C. M. Gutierrez    A. Hall    C. Hernández-Monteagudo    H. Hildebrandt    J. Hjorth    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    J. Kim    C. C. Kirkpatrick    S. Kruk    V. Le Brun    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    T. I. Liaudat    A. Loureiro    J. Macias-Perez    M. Magliocchetti    C. Mancini    R. Maoli    C. J. A. P. Martins    L. Maurin    M. Miluzio    P. Monaco    C. Moretti    G. Morgante    S. Nadathur    K. Naidoo    P. Natoli    A. Navarro-Alsina    S. Nesseris    D. Paoletti    F. Passalacqua    K. Paterson    L. Patrizii    A. Pisani    D. Potter    S. Quai    M. Radovich    P.-F. Rocci    G. Rodighiero    S. Sacquegna    M. Sahlén    D. B. Sanders    E. Sarpa    C. Scarlata    A. Schneider    D. Sciotti    E. Sellentin    F. Shankar    L. C. Smith    E. Soubrie    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    G. Verza    P. Vielzeuf    A. Viitanen    N. A. Walton    J. R. Weaver
@@
The slitless spectroscopy mode of the Near-Infrared Spectrometer and Photometer (NISP) on board the Euclid telescope has enabled efficient spectroscopy of objects within a large field of view. Nevertheless, the relatively low spectral resolution, overlapping spectra, and contamination pose challenges to source classification and redshift determination using the NISP spectra alone. In this work, we present a large and homogeneous sample of bright quasars identified from the Euclid Quick Data Release (Q1), constructed by combining high-purity candidate selections from Gaia and WISE with the new spectroscopic capabilities of Euclid. Through visual inspection of the Euclid spectra of these quasar candidates, we identify approximately 3500 quasars and determine reliable redshifts in the range of $0<z\lesssim 4.8$. Of these, 2686 are new spectroscopic identifications relative to existing public compilations. We generated the first Euclid composite spectrum of quasars covering rest-frame near-ultraviolet (NUV) to near-infrared (NIR) wavelengths without telluric lines, which will be pivotal to NIR quasar spectral analysis. We obtained an empirical spectroscopic depth of $J_{\scriptscriptstyle\rm E}\lesssim 21.5$ and $H_{\scriptscriptstyle\rm E}\lesssim 21.3$ at the sensitivity of the Wide Field Survey, beyond which the number of securely identified quasars declines sharply. Accordingly, the sample presented in this paper comprises spectroscopically confirmed quasars brighter than these limits. We analysed morphological parameters from the Visible Camera (VIS) using Sérsic and model-independent (CAS) metrics, and a deep-learning point spread function fraction to track nuclear dominance. The VIS morphologies show a clear redshift dependence: at low redshift ($z<0.5$), obvious host structures are common and a single Sérsic model fits about half of the sources; at intermediate redshift ($0.5<z<2$), the nuclear component dominates, with 90% of the Sérsic fits saturating at the upper index limit. In this intermediate redshift regime, $f_{{\scriptscriptstyle\rm PSF}}$ is available, and we use it as a more reliable compactness measure than the single-Sérsic and CAS parameters to quantify nuclear versus host emission. We also explore the novel Euclid NIR colour space and discuss the role of these quasars in refining active galactic nucleus selection techniques for future Euclid data releases.
@@ 1 Introduction
Powered by accreting supermassive black holes, quasars are among the most luminous and distant objects in the Universe. Quasars act as beacons that allow us to probe recent to early cosmic epochs and trace the large‐scale structure of the cosmos (e.g. Blanton et al. 2017; Neveux et al. 2020; Fan et al. 2023). Traditionally, quasar surveys have primarily relied on multi-colour selection techniques to isolate candidates, with subsequent slit or multi-fibre spectroscopy to confirm their nature and determine precise redshifts (e.g. Richards et al. 2002; Croom et al. 2004; Myers et al. 2015; Chaussidon et al. 2023). Ground-based quasar surveys have adopted slitless spectroscopy to find quasars in a cost-effective way. However, such slitless campaigns suffer from low spectral resolution, overlapping spectra, and contamination, often necessitating follow-up slit spectroscopy for confirmation (e.g. Schmidt et al. 1986; Osmer and Hewett 1991; Schneider et al. 1999).
@@
The advent of space-based observations and improvements in data reduction have substantially improved the quality of slitless spectroscopic data. For example, the Hubble Space Telescope’s (HST) Advanced Camera for Surveys (ACS) and Wide Field Camera 3 (WFC3) have produced high-quality grism data (e.g. Momcheva et al. 2016; Estrada-Carpenter et al. 2019) that are processed with dedicated pipelines (e.g. Kümmel et al. 2009; Brammer 2019). More recently, with the low background, high sensitivity, and high spatial resolution of the James Webb Space Telescope (JWST), the Near Infrared Camera (NIRCam; Rieke et al. 2005, 2023) and the Near Infrared Imager and Slitless Spectrograph (NIRISS; Willott et al. 2022; Doyon et al. 2023) on board JWST have enabled unprecedented studies of distant galaxies (e.g. Roberts-Borsani et al. 2022; Sun et al. 2023; Oesch et al. 2023; Meyer et al. 2024).
@@
The European Space Agency’s (ESA) Euclid mission (Euclid Collaboration: Mellier et al. 2025) is designed to probe the dark matter and dark energy of the Universe by studying weak lensing and galaxy clustering over approximately one third of the sky in both the optical and near-infrared using the Visible Camera (VIS; Euclid Collaboration: Cropper et al. 2025) and the Near-Infrared Spectrometer and Photometer (NISP; Euclid Collaboration: Jahnke et al. 2025). A key feature of Euclid is the slitless spectroscopic mode of NISP, which can simultaneously capture spectra of sources in a large field of view of 0.57 $\deg^{2}$. Three red grisms covering the same ${\rm RG}_{\scriptscriptstyle\rm E}$ band (1206–1892 nm) are adopted in the Euclid Wide Survey to provide spectra with different dispersion directions of 0°, 180°, and 270° with respect to the detector columns. Dispersed slitless images of the three grims are combined to disentangle overlapping spectra from multiple sources, and generate the final clean spectra. The red grims have a resolving power of ${{\cal R}_{{\rm{R}}{{\rm{G}}_{\rm{E}}}}}>480$ for a source with a 0.5 diameter (Euclid Collaboration: Jahnke et al. 2025; Euclid Collaboration: Mellier et al. 2025). This spectroscopic capability enables a census of bright quasars and enhances the efficiency of quasar discovery by leveraging the high sensitivity and spatial resolution from space. For example, Bañados et al. (2025) have recently discovered a $z=5.404$ quasar, EUCL J181530.01+652054.0, using NISP spectroscopy.
@@
Euclid Collaboration: Lusso et al. (2024) provide a detailed prediction of the NISP spectroscopic mode for active galactic nuclei using mock spectra. They demonstrate that redshift measurements are robust when the H $\alpha$ emission line is visible within the spectral coverage of ${\rm RG}_{\scriptscriptstyle\rm E}$ ($0.89<z<1.83$) at a line flux greater than $2\times 10^{-16}$ erg s-1 cm-2. Outside this redshift range, however, redshift measurements are inefficient due to a low signal-to-noise ratio (S/N) or a lack of prominent emission lines or both.
@@
Early investigations of AGNs in the first Euclid Quick Data Release (Q1; Euclid Collaboration: Aussel et al. 2025) already demonstrate the potential of Euclid for AGN science. For example, Euclid Collaboration: Matamoro Zatarain et al. (2025) present an AGN candidate catalogue with a total of 229 779 objects selected with multi-wavelength data, while Euclid Collaboration: Roster et al. (2025) identify Euclid counterparts to X-ray sources in the Deep Fields, most of which are AGNs, using catalogues from eROSITA (Merloni et al. 2024), XMM-Newton (Webb et al. 2020), and Chandra (Evans et al. 2024).
@@
In this work, we present a homogenous bright quasar sample identified with spectroscopic data of Q1, complementary to the efforts of Euclid Collaboration: Matamoro Zatarain et al. (2025) and Euclid Collaboration: Roster et al. (2025). We select quasar candidates from all-sky optical and mid-infrared databases, and identify the objects based on prominent emission lines in the Euclid slitless spectra. Our analysis demonstrates that slitless spectroscopy, as implemented in the Euclid mission, not only overcomes some of the limitations of colour-based quasar selection techniques, but also provides a valuable dataset for studying the large-scale structure of the Universe through quasar clustering and cross-correlation with galaxy and weak lensing maps (e.g. Myers et al. 2003; Pullen et al. 2015; Petter et al. 2023; Alonso et al. 2023). Spectral properties of these quasars obtained with multi-component fitting, including spectral indices, line widths, and black hole masses, will be reported in Euclid Collaboration: Calhau et al. (in preparation).
@@
The paper is organised as follows. Section 2 describes the sample selections of quasar candidates and the crossmatch to the Euclid spectral catalogue. Section 3 describes the identification and redshift determination procedure. We present our results in Sect. 4 and discuss their implications in Sect. 5. Finally, Sect. 6 concludes with a summary of our findings and prospects for future studies. All magnitudes are in the AB system (Oke and Gunn 1983) unless stated otherwise.